OpenStudy (anonymous):

Which two consecutive integers have the sum of -105

OpenStudy (anonymous):

Let the intergers be x and x+1 x + x + 1 = -105 2x = - 105 - 1 x = -53 The intergers are -53, and -52.

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

-53 -52